OpenStudy (anonymous):

3/2=(T+3)/(T-3) show your work

OpenStudy (anonymous):

hhhhhheeeeeeeelllllllllllllppppppppppppp!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (turingtest):

3/2=(T+3)/(T-3) (3/2)(T-3)=T+3 (3/2)T-9/2=T+3 T/2=15/2 T=15

OpenStudy (anonymous):

turing test is correct T=15 even his solution is correct